i dont understand this thank you verymu44ch



[(2+3) power of 2 -4.5}+7

Answer by fractalier(2101)

We use the order of operations, or PEMDAS, to do this properly...
First we work within the innermost parentheses and get
[(2 + 3)^2 - 4.5] + 7 =
[5^2 - 4.5] + 7 = then exponents
[25 - 4.5] + 7 = then within parentheses again
20.5 + 7 =
27.5